# RL for Reasoning by Adaptively Revealing Rationales

### v1

@@ 3 Experiments
In this section, we evaluate AdaBack on real-world reasoning tasks. In Sec. 2.2, we showed that AdaBack enables models to solve problems that are intractable under standard supervised fine-tuning, reinforcement learning, or their combination. Our goal here is to investigate whether AdaBack improves generalization and reasoning ability beyond what standard RL pipelines achieve on real-world reasoning tasks.
We train models using the GRPO algorithm [48] on base models of the LLaMA-3 family [23]. The standard RL recipe involves first performing SFT on rationales, followed by RL fine-tuning [25, 24]. However, this paradigm has recently been challenged by models trained solely with RL updates such as GRPO, without any SFT phase [16]. Accordingly, we compare four variants: GRPO on a base model, GRPO on an SFT model, AdaBack-GRPO on a base model, and AdaBack-GRPO on an SFT model. Further experimental details are provided in Appendix D.
@@
We evaluate on two standard mathematical reasoning benchmarks: MATH [27] and GSM8k [15]. In line with Mirzadeh et al. [38] and Li et al. [36], to assess generalization beyond pretraining exposure, we introduce two new variants of GSM8k:
@@
Table 1 summarizes performance across these tasks. We find that AdaBack consistently outperforms both GRPO and SFT+GRPO, particularly in out-of-distribution settings like Tensor-2 GSM8k. Notably, AdaBack applied directly to the base model often matches or exceeds the performance of the standard RL applied to the base and SFT-initialized models, suggesting that adaptive partial supervision on a base model may serve as a more effective prior than initializing from a SFT checkpoint.
@@
Method MATH GSM8k Base-7 GSM8k Tensor-2 GSM8k
1B 3B 1B 3B 1B 3B 1B 3B
Base+RL 6.4 15.0 7.9 63.7 4.8 4.9 0.0 0.0
SFT+RL 7.4 17.7 36.7 72.7 14.4 45.4 6.9 42.7
AdaBack 9.1 19.1 39.2 73.3 18.4 43.9 8.5 49.2
SFT+AdaBack 9.5 19.9 43.2 70.7 24.5 49.9 11.3 42.2
@@
Curriculum learning typically requires hand-crafted stages, scheduling heuristics, and careful tuning of hyperparameters—such as how long to train at each difficulty level and with what parameters, and when to transition to harder ones. In contrast, AdaBack performs an automatic per-sample curriculum by adjusting supervision based on a simple reward threshold. Each training point progresses at its own pace, without global stage definitions. Although our implementation uses a stochastic binary search over supervision ratios, this is not essential: linear search or other adaptive strategies could be substituted. The key principle is to allow each example to progress at its own pace, without requiring global curriculum design or extensive hyperparameter tuning. As seen in Figure 3 (left) average portions naturally decrease and rewards increase with AdaBack without any boilerplate curriculum scaffolding.
@@
Figure 3 shows training curves comparing AdaBack and standard RL across SFT and base model initializations. In both settings, AdaBack achieves a more stable reward trajectory and maintains a consistent gap between train and test performance—evidence of generalization to novel problems.
@@
Yue et al. [59] argue that RL fine-tuning reweights a model’s output distribution without expanding its effective reasoning capacity. To test this claim, we evaluate models using pass@k [^3] Pass@k is the probability that at least one out of $k$ model-generated outputs correctly solves a given problem.[13, 12], a metric that captures the breadth of plausible solutions. Figure 4 shows that AdaBack significantly improves pass@k over standard RL on both the base and SFT models, especially at large $k$. These gains are most pronounced when AdaBack is applied to the base model, again suggesting that it facilitates the discovery of new solution modes rather than just refining existing ones. If SFT-free RL with Adaback introduces novel capabilities, we expect pass@$k$ to increase even when base model coverage is low, which is what we show in Figure 4. In some cases, such as on the Tensor-2 GSM8k (see Table 1), AdaBack on the base model even outperforms AdaBack on the SFT-initialized model, suggesting that SFT may sometimes restrict the search space too early, limiting the benefits of exploration. The latter can be observed in Figure 4 as well.
Figure 4: Pass@k for LLama3-1B SFT-initialized models (left) and base models (right) on GSM8k. Adaback significantly improves performance at higher $k$ even without SFT suggesting it expands the solution distribution rather than reweighting known answers (contra [59]).
@@
On the MATH dataset with Llama 3.2 3B-Instruct, we observe no gains from AdaBack. As shown in Figure 5, both train and test rewards saturate quickly, with nearly all questions solved within a few hundred iterations. We observed a similar behavior on other models like the Qwen 2.5 family of models. This suggests that these models have likely encountered much of the dataset during pretraining, leaving little room for RL improvements or for Adaback to assist exploration. In such cases, AdaBack provides limited benefit—highlighting that its main strength lies in tasks where sparse reward or symbolic mismatch creates a significant learning barrier.
@@ 4 Related Work
Here, we provide a summary of the related work. We also provide a more extensive literature review in Appendix A.
@@
Several studies have highlighted the difficulty of applying reinforcement learning (RL) to structured reasoning tasks. Exploration remains a major bottleneck: reward signals are sparse, and correct reasoning chains are exponentially rare in the output space.
Havrilla et al. [26] empirically evaluated several RL strategies on reasoning benchmarks and found that models fail to discover solutions outside the support of the base SFT model. They tested curriculum-inspired techniques such as backtracking—starting the model partway through a solution and gradually shifting the start point earlier [43]—and prioritized level replay (PLR)—sampling more frequently from difficult problems [29]. Despite these interventions, performance gains were negligible: RL primarily amplified answers already assigned non-trivial probability by the pretrained model. These findings align with recent theoretical arguments by Yue et al. [59], who claim that RL fine-tuning reweighs existing reasoning paths but fails to induce fundamentally new capabilities.
@@
Curriculum learning was originally introduced to improve generalization and convergence by training models on simpler examples before harder ones [10]. In RL, backtracking [43] and RFCL [51] adopt this principle by initializing rollouts from later points in expert demonstrations and gradually shifting the start point backward.
In the context of language modeling, R3 [57] introduces a step-wise curriculum where the model is trained to complete reasoning chains from progressively earlier points. While motivated as a dynamic curriculum, R3 ultimately performs a static data augmentation: it slices demonstrations at all possible delimiter positions and treats each slice as an independent training sample. This approach depends critically on the existence of a consistent, domain-specific delimiter—such as newlines in GSM8k—to segment reasoning steps. In tasks without such natural structure, like our synthetic parity benchmark, this strategy is infeasible as shown in Fig. 2. Moreover, it requires handcrafted heuristics to identify step boundaries, limiting its generality.
@@
Several works have shown that allowing transformers to produce a chain-of-thought would increase their expressivity [18, 37]. Further, Abbe et al. [5] put forward the notion of globality degree of a task as a hardness measure and show that scratchpads can make learning more efficient by breaking the globality of a task. They also proposed inductive scratchpads which impose a Markovian structure over reasoning steps, improving length generalization. Gao et al. [20] proposed AbstRaL, showing improved robustness by training language models on mathematical abstraction of reasoning traces, instead of natural language chain-of-thoughts. Outside of reasoning and natural language scratch pads, unsupervised learning of intermediary symbolic sequences using straight-through gradient estimators [11] has been studied in [7, 44, 30].
5 Conclusion, Limitations, and Future Work
Curriculum learning is often seen as a training heuristic to improve convergence. In this work, we show that per-sample curriculum learning via adaptive partial supervision enables fundamentally new reasoning capabilities that are inaccessible to standard supervised fine-tuning or reinforcement learning alone. Our proposed method, AdaBack, succeeds on a synthetic parity task where both RL and SFT fail, offering a constructive separation result. On real-world datasets such as MATH [27], GSM8k [15], and its variants, AdaBack outperforms baseline RL and SFT+RL pipelines, often matching or exceeding standard RL applied on SFT-initialized models even when applied directly to a base model.
These results suggest that AdaBack not only improves sample efficiency, but also facilitates exploration of new solution modes, as evidenced by gains in pass@k. However, we also observe that AdaBack provides limited benefit when the base model has already memorized much of the task, as appears to be the case with modern reasoning-apt models and math datasets.
Appendix A Extended Related Work
@@
Reinforcement learning has emerged as a dominant paradigm for post-training language models on tasks involving sparse, verifiable reasoning signals [26]. Our reasoning tasks and settings are closest to STaR [60], Quiet-STaR [61], and R-STaR-MATH [24], which apply RL to improve mathematical and symbolic reasoning through selective training on verified rationales.
In parallel, alignment with human preferences has also leveraged RL techniques—most prominently Proximal Policy Optimization (PPO [47]) [8, 41]. Recent efforts have explored variants of REINFORCE-style updates and reward model bootstrapping, often using iterative fine-tuning pipelines. Interestingly, multiple works have reported that iterative filtering and fine-tuning on correct completions can match or exceed PPO performance in some domains [25, 6].
Two standard benchmarks for mathematical reasoning used in most of the just-mentioned work in RL for reasoning are GSM8k [15] and MATH [27] datasets. GSM8k consists of grade-school arithmetic problems requiring multi-step solutions, often accompanied by natural language justifications. MATH contains higher-difficulty competition-style math problems with structured step-by-step solutions. Nevertheless, RL approaches can be applied to any task for which verifiers or other forms of reward functions are available, including coding [13] and formal math [58].
@@
Several other works in reinforcement learning have explored curriculum strategies to overcome sparse reward challenges [39]. Florensa et al. [19] uses a reverse curriculum for training, where starting states become increasingly difficult during training. However, compared to Salimans and Chen [43], these starting states do not come from demonstrations. Backplay [42] demonstrated strong gains in environments like Pommerman and Atari by starting episodes near the goal state using a single demonstration and gradually moving the starting point backward, enabling the agent to outperform the suboptimal demonstrator. Prioritized Level Replay (PLR) [29] focuses training on the hardest levels in procedural environments like Procgen[14] by adaptively replaying levels where the agent performs poorly. Sukhbaatar et al. [50] proposed an automatic curriculum where a teacher agent proposes increasingly difficult tasks for a learner agent, leading to emergent complex curriculum strategies without any hand-designed task progression. This was extended in Sukhbaatar et al. [49] to learn goal embeddings and reusable low-level policies through self-play.

### v2

@@ 3 Experiments
In this section, we evaluate AdaBack on real-world reasoning tasks. In Section 2.2, we showed that AdaBack enables models to solve problems that are intractable under standard supervised fine-tuning, reinforcement learning, or their combination. Our goal here is to investigate whether AdaBack improves generalization and reasoning ability beyond what standard RL pipelines achieve on real-world reasoning tasks.
[…]
The standard RL recipe involves first performing SFT on rationales, followed by RL fine-tuning [Gulcehre et al., 2023, Guan et al., 2025]. However, this paradigm has recently been challenged by models trained solely with RL updates such as GRPO, without any SFT phase [DeepSeek-AI, 2025]. Accordingly, we compare four variants: GRPO on a base model, GRPO on an SFT model, AdaBack-GRPO on a base model, and AdaBack-GRPO on an SFT model. Further experimental details are provided in Appendix D.
@@
We evaluate on three standard mathematical reasoning benchmarks: DeepScaleR [Luo et al., 2025], MATH [Hendrycks et al., 2021] and GSM8k [Cobbe et al., 2021]. In line with Mirzadeh et al. [2024] and Li et al. [2024], to assess generalization beyond pretraining exposure, we introduce two new variants of GSM8k:
@@
Table 1 summarizes performance across these tasks. We find that AdaBack consistently outperforms both GRPO and SFT+GRPO, particularly in out of pretraining distribution settings like Tensor-2 GSM8k. Notably, AdaBack applied directly to the base model often matches or exceeds the performance of the standard RL applied to the base and SFT-initialized models, suggesting that adaptive partial supervision on a base model may serve as a more effective prior than initializing from a SFT checkpoint. Interestingly, AdaBack applied on the base model may even outperform AdaBack applied on the SFT model as in the Tensor-2 experiment. We elaborate on this phenomenon further in the text.
@@
Method DeepScaleR MATH GSM8k Base-7 GSM8k Tensor-2 GSM8k
1B 3B 1B 3B 1B 3B 1B 3B 1B 3B
Base+RL 6.8 6.6 6.4 15.0 7.9 63.7 4.8 4.9 0.0 0.0
SFT+RL 7.1 9.1 7.4 17.7 36.7 72.7 14.4 45.4 6.9 42.7
AdaBack 9.0 10.6 9.1 19.1 39.2 73.3 18.4 43.9 8.5 49.2
SFT+AdaBack 9.5 12.5 9.5 19.9 43.2 70.7 24.5 49.9 11.3 42.2
@@
Curriculum learning typically requires hand-crafted stages, scheduling heuristics, and careful tuning of hyperparameters—such as how long to train at each difficulty level and with what parameters, and when to transition to harder ones. In contrast, AdaBack performs an automatic per-sample curriculum by adjusting supervision based on a simple reward threshold. Each training point progresses at its own pace, without global stage definitions. Although our implementation uses a stochastic binary search over supervision ratios, this is not essential: linear search or other adaptive strategies could be substituted. The key principle is to allow each example to progress at its own pace, without requiring global curriculum design or extensive hyperparameter tuning. As seen in Figure 3 (left) average portions naturally decrease and rewards increase with AdaBack without any boilerplate curriculum scaffolding. $\tau$ is the only hyperparameter needed to be set for AdaBack, and training is not sensitive to it. We have elaborated the choice of $\tau$ in Section C.
@@
Figure 3 shows training curves comparing AdaBack and standard RL across SFT and base model initializations. In both settings, AdaBack achieves better rewards both at training and test compared to standard GRPO training. We provide additional training figures in Appendix F.
@@
Yue et al. [2025] argue that RL fine-tuning reweights a model’s output distribution without expanding its effective reasoning capacity. To test this claim, we evaluate models using pass@k [^3] Pass@k is the probability that at least one out of $k$ model-generated outputs correctly solves a given problem.Chen et al. [2021], Brown et al. [2024], a metric that captures the breadth of plausible solutions. Figure 4 shows that AdaBack significantly improves pass@k over standard RL on both the base and SFT models, especially at large $k$. These gains are most pronounced when AdaBack is applied to the base model, again suggesting that it facilitates the discovery of new solution modes rather than just refining existing ones. If SFT-free RL with AdaBack introduces novel capabilities, we expect pass@$k$ to increase even when base model coverage is low, which is what we show in Figure 4. In some cases, such as on the Tensor-2 GSM8k (see Table 1), AdaBack on the base model even outperforms AdaBack on the SFT-initialized model, suggesting that SFT may sometimes restrict the search space too early, limiting the benefits of exploration. [^4] In an extreme case, the model may just memorize the answers during the SFT phase, removing hope for any exploration. The latter can be observed in Figure 4 as well.
Figure 4: Pass@k for Llama3-1B SFT-initialized models (left) and base models (right) on GSM8k. AdaBack keeps a significant gap compared to standard RL and improves performance at higher $k$ even without SFT suggesting it expands the solution distribution rather than reweighting known answers (contra Yue et al. [2025]).
@@
On the MATH dataset with Llama 3.2 3B-Instruct, we observe no gains from AdaBack. As shown in Figure 10, both train and test rewards saturate quickly, with nearly all questions solved within a few hundred iterations. In Figure 11 we illustrate the training and test reward/accuracy trajectories for Qwen2.5-1.5B (both base and instruct) [Yang et al., 2024] on the GSM8k dataset, where we observed a similar behavior. We observed the same phenomena in the case of GSM8k for the 3B model in Table 1. This happens when RL reaches perfect training reward within a few iterations during training, i.e., when the dataset is too simple for the model. This suggests that these models have likely encountered much of the dataset during pretraining, leaving little room for RL improvements or for AdaBack to assist exploration. In such cases, AdaBack provides limited benefit—highlighting that its main strength lies in tasks where sparse reward or symbolic mismatch creates a significant learning barrier.
Comparison with R3 on GSM8k, MATH, and DeepScaleR
Table 2 reports a direct comparison between AdaBack and R3 on three datasets of increasing difficulty—GSM8k < MATH < DeepScaleR [Luo et al., 2025]—using Llama-3 1B and Llama-3 3B base models. Across all settings, AdaBack and R3 rely on the same ground-truth solutions and reward functions; the only difference is the curriculum mechanism. This isolates the effect of adaptive per-sample supervision from the mixture approach of R3.
Overall, we observe a clear trend aligned with dataset difficulty. On DeepScaleR, AdaBack outperforms R3 for both 1B and 3B models. On MATH, AdaBack has an edge for the 1B model, though this edge diminishes for the larger 3B model. On GSM8k, AdaBack performs slightly below R3, though the absolute gap remains small relative to the inherent variance of RL post-training.
We believe two structural factors underlie this pattern:
1. Segmentation brittleness on complex reasoning traces. MATH and DeepScaleR contain long, multi-line reasoning steps, LaTeX/MathJax blocks, and irregular formatting. These properties make R3’s slicing into fixed “stages” brittle and sensitive to heuristic delimiter choices, whereas AdaBack eliminates the need for any segmentation hyperparameters by using reward-driven adaptivity. On the other hand, for GSM8k, different steps of the solution can be easily separated using the newline delimiter. The latter can explain he superior performance of R3 and suggests that slicing solutions based on individual reasoning steps (as in R3) may work better than slicing the solution at random points (as done by AdaBack) when clean segmentation is possible. However, as discussed earlier, separating reasoning steps is not always feasible (e.g., in MATH and DeepScaleR).
2. Per-sample adaptivity becomes increasingly important on heterogeneous, long-horizon tasks. Harder datasets such as MATH and DeepScaleR contain many questions of widely varying structure and solution length. R3 differs from AdaBack in two aspects:
  1. 1.
    It slices all solutions into an equal number of pieces (e.g., 5), which is set globally for the whole dataset. Such global slicing does not account for variations in question difficulty and cannot adapt to intra-dataset heterogeneity. AdaBack’s per-sample curriculum, which adjusts the supervision length independently for each example, is particularly advantageous in these settings and supports more stable progression on difficult problems.
  2. 2.
    Instead of imposing a curriculum over the data, R3 creates a mixture from all questions and sliced solutions. This contrasts with AdaBack’s easy-to-hard curriculum. The use of a curriculum becomes increasingly important as the difficulty of the training set increases. In particular, Abbe et al. [2023b] provide theoretical results comparing curriculum learning and data mixtures: curriculum learning is beneficial when easy samples within a distribution are rare; conversely, if such samples are prevalent, standard training on the mixture may be more efficient than a curriculum.
To further support our argument, we refer the reader back to the chain-of-parities experiment presented in Figure 2, where it is shown that as the number of (non-trivial) reasoning steps in a task increases, the AdaBack approach can significantly outperform R3 (making training $20\times$ more efficient in Figure 2). The real-world datasets used in our paper still contain only a few reasoning steps. We believe AdaBack will better demonstrate its advantages on longer-horizon reasoning tasks (e.g., math Olympiad problems). However, we point out that experimenting in such regimes presents major challenges regarding data acquisition and large-scale compute.
Table 2: Comparison of R3 and AdaBack on GSM8k, MATH, and DeepScaleR
Method GSM8k MATH DeepScaleR
1B (base) 3B (base) 1B (base) 3B (base) 1B (base) 3B (base)
R3 41.5 74.2 7.8 19.2 6.6 9.6
AdaBack 39.2 73.3 9.1 19.1 9.0 10.6
@@ 4 Related Work
We also provide a more extensive literature review in Appendix A.
@@
Several studies have highlighted the difficulty of applying reinforcement learning (RL) to structured reasoning tasks. Exploration remains a major bottleneck: reward signals are sparse, and correct reasoning chains are exponentially rare in the output space. Havrilla et al. [2024] empirically evaluated several RL strategies on reasoning benchmarks and found that models fail to discover solutions outside the support of the base SFT model. They tested curriculum-inspired techniques such as backtracking—starting the model partway through a solution and gradually shifting the start point earlier Salimans and Chen [2018]—and prioritized level replay (PLR)—sampling more frequently from difficult problems Jiang et al. [2020]. Despite these interventions, performance gains were negligible: RL primarily amplified answers already assigned non-trivial probability by the pretrained model. These findings align with recent theoretical arguments by Yue et al. [2025], who claim that RL fine-tuning reweighs existing reasoning paths but fails to induce fundamentally new capabilities.
@@
Curriculum learning was originally introduced to improve generalization and convergence by training models on simpler examples before harder ones [Bengio et al., 2009]. In RL, backtracking [Salimans and Chen, 2018] and RFCL Tao et al. [2024] adopt this principle by initializing rollouts from later points in expert demonstrations and gradually shifting the start point backward.
In language modeling, R3 [Xi et al., 2024] introduces a step-wise curriculum where the model is trained to complete reasoning chains from progressively earlier points. However, despite being motivated as a dynamic curriculum, R3 effectively performs static data augmentation: demonstrations are sliced at delimiter positions (e.g., newlines in GSM8k), and each slice is treated as an independent training sample. In datasets without consistent delimiters, R3 falls back on uniformly partitioning each solution into a fixed number of segments (e.g., five), regardless of problem difficulty or solution length, and then mixing all segments across the dataset. This reliance on global heuristics limits scalability and applicability to domains without clear step boundaries, such as MATH [Hendrycks et al., 2021] or our synthetic parity benchmark. In particular, for the parity task we show that R3 has significantly worse learning speed and generalization in Figure 2. This is consistent with the theoretical results of Cornacchia and Mossel [2023], Abbe et al. [2023b] where it is proven that applying a curriculum over distributions can significantly improve the speed of convergence over simply mixing the distributions.
@@
5 Conclusion, Limitations, and Future Work
Curriculum learning is often seen as a training heuristic to improve convergence. In this work, we show that per-sample curriculum learning via adaptive partial supervision enables fundamentally new reasoning capabilities that are inaccessible to standard supervised fine-tuning or reinforcement learning alone. Our proposed method, AdaBack, succeeds on a synthetic parity task where both RL and SFT fail, offering a constructive separation result. On real-world datasets such as MATH [Hendrycks et al., 2021], GSM8k [Cobbe et al., 2021], and its variants, AdaBack outperforms baseline RL and SFT+RL pipelines, often matching or exceeding standard RL applied on SFT-initialized models even when applied directly to a base model.
These results suggest that AdaBack not only improves sample efficiency, but also facilitates exploration of new solution modes, as evidenced by consistent gains in pass@k. However, we also observe that AdaBack provides limited benefit when the base model has already memorized much of the task, as appears to be the case with modern reasoning-apt models and math datasets.
Appendix A Extended Related Work
@@
Several works have shown that allowing transformers to produce a chain-of-thought would increase their expressivity [Feng et al., 2024, Merrill and Sabharwal, 2024]. Further, Abbe et al. [2024] put forward the notion of globality degree of a task as a hardness measure and show that scratchpads can make learning more efficient by breaking the globality of a task. They also proposed inductive scratchpads which impose a Markovian structure over reasoning steps, improving length generalization. Gao et al. [2025] proposed AbstRaL, showing improved robustness by training language models on mathematical abstraction of reasoning traces, instead of natural language chain-of-thoughts. Outside of reasoning and natural language scratch pads, unsupervised learning of intermediary symbolic sequences using straight-through gradient estimators [Bengio et al., 2013] has been studied in [Amani et al., 2024, Sánchez et al., 2023, Kaiser and Bengio, 2018].
@@
Reinforcement learning has emerged as a dominant paradigm for post-training language models on tasks involving sparse, verifiable reasoning signals Havrilla et al. [2024]. Our reasoning tasks and settings are closest to STaR [Zelikman et al., 2022], Quiet-STaR [Zelikman et al., 2024], and R-STaR-MATH [Guan et al., 2025], which apply RL to improve mathematical and symbolic reasoning through selective training on verified rationales.
In parallel, alignment with human preferences has also leveraged RL techniques—most prominently Proximal Policy Optimization (PPO [Schulman et al., 2017]) [Bai et al., 2022, Pang et al., 2024b]. Recent efforts have explored variants of REINFORCE-style updates and reward model bootstrapping, often using iterative fine-tuning pipelines. Interestingly, multiple works have reported that iterative filtering and fine-tuning on correct completions can match or exceed PPO performance in some domains [Gulcehre et al., 2023, Ahmadian et al., 2024].
Two standard benchmarks for mathematical reasoning used in most of the just-mentioned work in RL for reasoning are GSM8k [Cobbe et al., 2021] and MATH [Hendrycks et al., 2021] datasets. GSM8k consists of grade-school arithmetic problems requiring multi-step solutions, often accompanied by natural language justifications. MATH contains higher-difficulty competition-style math problems with structured step-by-step solutions. Nevertheless, RL approaches can be applied to any task for which verifiers or other forms of reward functions are available, including coding [Chen et al., 2021] and formal math [Yang et al., 2023].
Alternatively to RL, some natural language questions can be translated into formal logic and answered using automated theorem provers—for instance, Olausson et al. [2023] maps questions into first-order logic using an LLM, then delegates inference to a symbolic prover. AbstRaL also solves GSM8k-style problems by first converting them into abstract symbolic equations and then using a solver [Gao et al., 2025].
@@
Several other works in reinforcement learning have explored curriculum strategies to overcome sparse reward challenges [Narvekar et al., 2020]. Florensa et al. [2017] uses a reverse curriculum for training, where starting states become increasingly difficult during training. However, compared to Salimans and Chen [2018], these starting states do not come from demonstrations. Backplay [Resnick et al., 2022] demonstrated strong gains in environments like Pommerman and Atari by starting episodes near the goal state using a single demonstration and gradually moving the starting point backward, enabling the agent to outperform the suboptimal demonstrator. Prioritized Level Replay (PLR) [Jiang et al., 2020] focuses training on the hardest levels in procedural environments like Procgen[Cobbe et al., 2019] by adaptively replaying levels where the agent performs poorly. Sukhbaatar et al. [2018b] proposed an automatic curriculum where a teacher agent proposes increasingly difficult tasks for a learner agent, leading to emergent complex curriculum strategies without any hand-designed task progression. This was extended in Sukhbaatar et al. [2018a] to learn goal embeddings and reusable low-level policies through self-play.
Another subset of works in curriculum learning [Chen et al., 2025, Shi et al., 2025, Bae et al., 2025] primarily adjust dataset/batch-level curricula, i.e., selecting which questions to include in each RL batch. By contrast, AdaBack introduces a per-sample curriculum, adaptively revealing partial rationales while sampling questions uniformly. Thus these approaches are complementary: AdaBack can be combined with batch-level curricula but unlike these works which do not benefit from chain-of-thought (CoT) demonstrations, AdaBack uses CoT supervised data to further improve efficiency.
Comparison with State-of-the-art Mathematical Reasoning Models
Although recent leader-board results show strong performance, they are not directly comparable to our work because they typically rely on much larger models, additional math-heavy pretraining datasets, or complex prompting and inference-time searching strategies S. Xue, Z. Huang, J. Liu, X. Lin, Y. Ning, B. Jin, X. Li, and Q. Liu (2024), R. Y. Pang, W. Yuan, K. Cho, H. He, S. Sukhbaatar, and J. E. Weston (2024a), 61. In contrast, our focus is on isolating the contribution of AdaBack itself by comparing it fairly against standard RL baselines under controlled settings. Importantly, AdaBack is not an alternative to these methods but a complementary training enhancement. It can be combined with larger models, post-training strategies, or prompting techniques like chain-of-thought prompting and majority voting Xue et al. [2024] without overhead. This makes AdaBack a foundational component that strengthens models from the ground up and can integrate seamlessly into more complex state-of-the-art pipelines Pang et al. [2024a].